OpenStudy (anonymous):

Use the trapezoid rule to approximate the area of the pool formed by a water main break. The measurements were taken 30cm apart.

OpenStudy (anonymous):

Composite Trapezoid Rule: int(f(x),x=a..b)=((b-a)/(2*n)) * (f(a)+2*sum(f(xj),j=1..n-1)+f(b)) So for you, int(f(x),x=a..b)=15*(f(a)+2*f(a+30)+2*f(a+60)+...+f(b)) I assume your first and lest depth measurements are zero, so int(f(x),x=a..b)=30*(sum of all depth measurements)